OpenStudy (anonymous):

Determine the equation of the line, in slope-intercept form, that will get your spacecraft from (1,2) to (5,5).

OpenStudy (anonymous):

So far I did the work --- m= 5-2/5-1= ¾. Then you put in into point-slope form, y-5= ¾ (x-5). Next multiply 4 on both sides, (4)y – (4)5= ¾ (4) (x-5). 4y- 20 = 3(x-5). Then simplify the right side, 4y – 20 = 3x -15. Then add 20 on both sides, 4y = 3x + 5. I'm not sure if my answer is 4y = 3x +5 or if the y part in slope-intercept form should be alone?

OpenStudy (anonymous):

@amistre64 can you help me please?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

i'm sorry to bother you guys

OpenStudy (anonymous):

i guess the solution is 4y=3x+5

OpenStudy (nurali):

\[m=\frac{ 5-2 }{ 5-1 }\] \[m=\frac{ 3 }{ 4 }\] \[y-y _{1}=m(x-x _{1})\] \[y-2=3/4(x-1)\] \[4y-8=3x-3\] \[4y=3x+5\]

OpenStudy (anonymous):

@nurali:ya i did the same way

OpenStudy (anonymous):

slope intercept form is y=mx+c then why the 4y is there..? PS- divide 4 on both sides

OpenStudy (anonymous):

ok so for y=mx +b the answer would be 4y = 3x +5

OpenStudy (anonymous):

ya...if u want in that format..then ur answer will be y=(3/4)x+(5/4)

OpenStudy (anonymous):

@divu.mkr how do you divide 4 on both sides if you got an x on one side and a y on another? I think 4y = 3x +5 is right. if not i'll write the right one later when I find out :)

OpenStudy (anonymous):

the slope-intercept is y=mx+b .. you need to divide 4 on the both sides to satisfy the slope-intercept .. :)

OpenStudy (anonymous):

ya right...so ultimately the answer will be y=(3/4)x+(5/4)..

OpenStudy (anonymous):

@susanka is right.. :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i'm going to double check because it could go any way :) but you guys could be right :)

OpenStudy (anonymous):

okay, girl . :)

OpenStudy (anonymous):

I believe you gusy areright because I did a different method for solving this problem so naturally it should be y =3/4x + 5/4 :)

OpenStudy (anonymous):

*guys are right

OpenStudy (anonymous):

that's good. :) ^_^

OpenStudy (anonymous):

dear slope intercept is y=mx+c not ay=bx+c @issy987

OpenStudy (anonymous):

Hey sorry for bothering again but I got one more question... if the answer is y = 3/4x +5/4 shouldn't the b intercept be a whole number?

OpenStudy (amistre64):

(1 ,2) to (5,5) -1-2 -1-2 ------------ 0,0 4,3 x = 4t + 1 ; solve for t, (x-1)/4 y = 3t + 2 ; plug in for t y = 3(x-1)/4 + 2 y = 3x/4 -3/4 + 8/4 y = 3x/4 + 5/4

OpenStudy (anonymous):

not necessary @issy987

OpenStudy (amistre64):

the y intercept is whatever real number the line crosses the y axis at.